A 50 kg shell is fired from a cannon on Earth's surface to a maximum height of 400 m. It then falls to a height of 100 m. What is the loss of potential energy as it falls?




1
Expert's answer
2021-02-05T02:16:43-0500

We can find the loss in potential energy as follows:


"\Delta PE=mg(h_f-h_i)=50\ kg\cdot9.8\cdot(100\ m-400\ m)=-147000\ J."